# Variations of metric that preserve a Riemannian submersion and geometry of its fibers

### v1

@@ 3.2 Preserving extrinsic geometry of fibers
for all $X,Y\in\mathfrak{X}_{\ker\pi_{*}}$, i.e., for all $x\in U$ the restriction of $\xi$ to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a Killing vector field. Since $\xi\neq 0$ at some $y\in\pi^{-1}(U)$, let $\rho$ be a smooth function compactly supported in $U$ and such that $\rho(\pi(y))>0$, then $(\rho\circ\pi)\cdot\xi$ is a smooth vertical vector field on $M$, whose restriction to every fiber of $\pi$ is a Killing field.
On the other hand, suppose that on $(M,g_{0})$ there exists a bounded non-trivial vertical vector field $\xi$, whose restriction to every fiber is Killing. Let $\{W_{n+1},\ldots,W_{n+p}\}$ be linearly independent vector fields on an open set $U\subset N$, such that $\xi\neq 0$ at some $x\in\pi^{-1}(U)$. We define a variation ${\tilde{g}}_{t}$ on $\pi^{-1}(U)$ by Corollary 3.6, setting $V_{n+1}=\xi$ and $V_{n+2}=\ldots=V_{n+p}=0$, and then use a non-zero, compactly supported in $U$ function $\rho$, such that $0\leq\rho\leq 1$ and $\rho(\pi(x))>0$, to define a variation $g_{t}$ on $M$ by $\partial_{t}g_{t}=(\rho\circ\pi)\cdot\partial_{t}{\tilde{g}}_{t}$. ∎
@@
To define a variation that preserves a Riemannian submersion with totally geodesic fibers, it is sufficient to assume that there exists a bounded Killing vector field $\xi$ on a fiber $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for one $x\in N$, because all fibers of a Riemannian submersion with totally geodesic fibers are isometric [8] and we can extend such $\xi$ to a vertical field on $(M,g_{0})$, whose restriction to every fiber is a Killing vector field.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with minimal fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with minimal fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is divergence-free for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is divergence-free, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with totally umbilical fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally umbilical fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a conformal Killing field for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is conformal Killing, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
$2P_{{\cal H}}^{t}(\nabla^{t}_{X}Y-\frac{1}{n}g_{t}(X,Y)H_{t})$ $=$ $\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{X}V_{i},Y)+g_{0}(\nabla^{0}_{Y}V_{i},X)$ (28)
$-\frac{2}{n}g_{t}(X,Y)\operatorname{div}^{0}_{\ker\pi_{*}}V_{i}\big{)}P_{{\cal H}}^{t}\pi^{*}W_{i}.$
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with minimal fibers. Then for every open set $U\subset M$ there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ such that $g_{t}=g_{0}$ on $M\setminus U$ and $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with minimal fibers for all $t\in(-\epsilon,\epsilon)$.
@@
Let $x\in\pi(U)$ and let ${\cal P}_{x}$ be the differential operator $\operatorname{div}^{0}_{\ker\pi_{*}}$ acting on the space $\mathfrak{X}_{{\cal F}_{x}}$ of vector fields on the fiber ${\cal F}_{x}$. Then ${\cal P}_{x}$ is underdetermined elliptic, i.e., its symbol is surjective and non-injective [2]. Hence, ${\cal P}_{x}$ has infinite-dimensional kernel ${\cal K}_{x}$, which includes an infinite set of compactly supported vector fields on ${\cal F}_{x}$ [5]. Let $\{V_{n+1},\ldots,V_{n+p}\}$ be such vector fields. We can assume that $\pi(U)$ is a normal coordinate neighbourhood of $x$; for every $y\in\pi(U)$ let $\phi_{x,y}:{\cal F}_{x}\rightarrow{\cal F}_{y}$ be the map induced by the flow of the $g_{0}$-horizontal lift of the vector field tangent to the geodesic from $x$ to $y$. Since the fibers are minimal, the maps $\phi_{x,y}$ preserve the volume form induced on the fibers (see e.g., [10, Chapter XV, Theorem 6.6]). Hence, extending $\{V_{n+1},\ldots,V_{n+p}\}$ to $U$ by maps $\phi_{x,y*}$, we obtain a set of vertical vector fields, such that the restriction to every fiber in $U$ of each of them is divergence-free and compactly supported on the fiber. Let $\rho$ be a smooth compactly supported function on $\pi(U)$. Multiplying each $V_{i}$ by $\rho\circ\pi$ we obtain a set of compactly supported, vertical vector fields in kernel of $\operatorname{div}^{0}_{\ker\pi_{*}}$, that by Corollary 3.6 defines a family $g_{t}$ with the required properties. ∎
@@ 4 Variational problems
Let $\pi:(M,g_{t})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\operatorname{vol}_{g_{t}}$ be the volume form on $(M,g_{t})$. For $(0,2)$-tensor fields $Q,S$ on $M$, let
@@
where $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{\ker\pi_{*}}$ is a set of local $g_{0}$-orthonormal vector fields.
Let $J_{h},J_{H}:\operatorname{Riem}(M,\ker\pi_{*},g_{0})\rightarrow\mathbb{R}$ be functionals defined by formulas
@@
where $\|h_{t}\|^{2}=\sum\nolimits_{a=1}^{n}g_{t}(h_{t}(E_{a},E_{b}),h_{t}(E_{a},E_{b}))$ for a local $g_{t}$-orthonormal frame $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{\ker\pi_{*}}$. We say that a metric $g_{0}$ is critical for $J_{h}$ (resp. $J_{H}$) with respect to variations preserving the Riemannian submersion $\pi$, if for all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$, such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$, we have $\partial_{t}J_{h}(g_{t})|_{t=0}=0$ (resp. $\partial_{t}J_{H}(g_{t})|_{t=0}=0$).
@@
We note that all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserving the Riemannian submersion $\pi$ also preserve the volume form $\operatorname{vol}_{g_{0}}$, since $\partial_{t}\operatorname{vol}_{g_{t}}=\frac{1}{2}\operatorname{Tr\,}B_{t}^{\sharp}\operatorname{vol}_{g_{t}}$ [3]. Indeed, let $\{E_{a},{\cal E}_{i}(t)\}$ be an adapted frame as in Lemma 3.8, then by (2) and (3) we have
@@ 4.1 Critical metrics
We will use notation from Section 2, especially its last paragraph. In particular, we have $P_{{\cal H}}^{0}(\delta_{0}h_{0})(X)=P_{{\cal H}}^{0}\sum\nolimits_{a=1}^{n}(\nabla^{0}_{E_{a}}h_{0})(E_{a},X)$ for all $X\in\mathfrak{X}_{\ker\pi_{*}}$ and any $g_{0}$-orthonormal local frame $\{E_{1},\ldots,E_{n}\}\subset\mathfrak{X}_{\ker\pi_{*}}$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Then $g_{0}$ is critical for $J_{h}$, with respect to variations preserving the Riemannian submersion $\pi$, if and only if $P_{{\cal H}}^{0}(\delta_{0}h_{0})=0$.
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a one-parameter family of Riemannian metrics on $M$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$.
Let $\{U_{\alpha},\alpha\in\{1,\ldots,\Lambda\}\}$ be a finite open cover of $N$, such that on every $U_{\alpha}$ there exists a frame of linearly independent vector fields $\{W_{n+1}^{\alpha},\ldots,W_{n+p}^{\alpha}\}$ and on $\pi^{-1}(U_{\alpha})$ we have $\operatorname{vol}_{g_{0}}=\pi^{*}(\operatorname{vol}_{g_{N}}(x))\otimes\operatorname{vol}_{{\cal F}_{x},g_{0}}$ for all $x\in U_{\alpha}$ [10], where $\operatorname{vol}_{{\cal F}_{x},g_{0}}$ is the volume form of ${\cal F}_{x}$ induced from $(M,g_{0})$. Then, there exist vertical vector fields $\{V_{n+1}^{\alpha},\ldots,V_{n+p}^{\alpha}\}$ on $\pi^{-1}(U_{\alpha})$ such that
$B_{t}^{\sharp}X=\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i}^{\alpha},X)P_{{\cal H}}^{t}\pi^{*}W_{i}^{\alpha}$
for all vertical vector fields $X$ on $\pi^{-1}(U_{\alpha})$. Let $z\in\pi^{-1}(U_{\alpha})$, then for every set of local $g_{0}$-orthonormal vector fields $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{\ker\pi_{*}}$ in a neighbourhood of $z$ we have, by (3.2), at the point $z$:
@@
$=2g_{t}((\partial_{t}\nabla^{t})_{E_{a}}E_{b},h_{t}(E_{a},E_{b}))$
$=\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{E_{a}}V^{\alpha}_{i},E_{b})+g_{0}(\nabla^{0}_{E_{b}}V^{\alpha}_{i},E_{a})\big{)}g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i}^{\alpha},h_{t}(E_{a},E_{b}))$
$=\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V^{\alpha}_{i}}g_{0},h^{i,\alpha}_{t}\rangle_{{\cal V}},$ (31)
@@
$=\sum\nolimits_{\alpha=1}^{\Lambda}\sum\nolimits_{i=n+1}^{n+p}\int_{\pi^{-1}(U_{\alpha})}f_{\alpha}\langle\mathcal{L}_{V_{i}^{\alpha}}g_{0},h^{i,\alpha}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}$
$=\sum\nolimits_{\alpha=1}^{\Lambda}\sum\nolimits_{i=n+1}^{n+p}\int_{U_{\alpha}}f_{\alpha}\big{(}\int_{{\cal F}_{x}}\langle\mathcal{L}_{V_{i}}g_{0},h^{i,\alpha}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{{\cal F}_{x},g_{0}}\big{)}\pi^{*}(\operatorname{vol}_{g_{N}}(x)).$ (32)
We will prove that $\partial_{t}J_{h}(g_{t})|_{t=0}=0$ for all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserving the Riemannian submersion $\pi$ if and only if for all $x\in N$, all $V\in\mathfrak{X}_{{\cal F}_{x}}$ and all $W\in T_{x}N$ we have
@@
We then set $V_{n+2}=\ldots=V_{n+p}=0$ and define a variation ${\tilde{g}}_{t}$ on $\pi^{-1}(U)$ as in Corollary 3.6. Using a non-zero, compactly supported in $U$ function $\rho$, such that $0\leq\rho\leq 1$ we define a variation $g_{t}$ on $M$ by $\partial_{t}g_{t}=\rho\cdot\partial_{t}{\tilde{g}}_{t}$, for which $\partial_{t}J_{h}(g_{t})|_{t=0}\neq 0$.
@@
We note that for all $X\in\mathfrak{X}_{\ker\pi_{*}}$ and for a local $g_{0}$-orthonormal adapted frame $\{E_{a},{\cal E}_{i}\}$, we can write the Codazzi formula [15] in the following form:
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. If the fibers of $\pi$ are totally umbilical on $(M,g_{0})$, then $g_{0}$ is critical for $J_{h}$, with respect to variations preserving the Riemannian submersion $\pi$, if and only if all fibers of $\pi$ are of parallel mean curvature on $(M,g_{0})$, i.e., $P_{{\cal H}}^{0}\nabla^{0}_{X}H_{0}=0$ for all $X\in\mathfrak{X}_{\ker\pi_{*}}$.
@@
For totally umbilical fibers $P_{{\cal H}}^{0}(\delta_{0}h_{0})(X)=\frac{1}{n}P_{{\cal H}}^{0}\nabla^{0}_{X}H_{0}$ for all $X\in\mathfrak{X}_{\ker\pi_{*}}$. ∎
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Then $g_{0}$ is critical for $J_{H}$, with respect to variations preserving the Riemannian submersion $\pi$, if and only if $H_{0}$ is projectable.
@@
$\int_{{\cal F}_{x}}g_{0}(H_{0},\pi^{*}W)\cdot(\operatorname{div}^{0}_{\ker\pi_{*}}V)\;\operatorname{vol}_{{\cal F}_{x},g_{0}}=0.$ (34)
@@
It follows that $g_{0}$ is critical for (30) with respect to variations preserving the Riemannian submersion $\pi$ if and only if $P_{{\cal V}}^{0}\nabla^{0}g_{0}(H_{0},\pi^{*}W)=0$ for all $W\in\mathfrak{X}_{N}$. Writing $H_{0}=\sum\nolimits_{i=n+1}^{n+p}g_{0}(H_{0},\pi^{*}W_{i})\pi^{*}W_{i}$ for a local $g_{N}$-orthonormal basis $\{W_{n+1},\ldots,W_{n+p}\}$, we find that this condition is equivalent to $H_{0}$ being projectable, as for all $V\in\mathfrak{\ker\pi_{*}}$
@@
We note that for a fixed $x\in N$, we can define functionals by (29) and (30), with $M$ in their definitions replaced by the fiber ${\cal F}_{x}$. The critical metrics then satisfy the same conditions as those given in Theorems 4.2 and 4.5, only on the fiber ${\cal F}_{x}$ rather than everywhere on $M$.

### v2

@@ 3.2 Preserving extrinsic geometry of fibers
for all $X,Y\in\mathfrak{X}_{{\cal V}}$, i.e., for all $x\in U$ the restriction of $\xi$ to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a Killing vector field. Since $\xi\neq 0$ at some $y\in\pi^{-1}(U)$, let $\rho$ be a smooth function compactly supported in $U$ and such that $\rho(\pi(y))>0$, then $(\rho\circ\pi)\cdot\xi$ is a smooth vertical vector field on $M$, whose restriction to every fiber of $\pi$ is a Killing field.
On the other hand, suppose that on $(M,g_{0})$ there exists a bounded non-trivial vertical vector field $\xi$, whose restriction to every fiber is Killing. Let $\{W_{n+1},\ldots,W_{n+p}\}$ be linearly independent vector fields on an open set $U\subset N$, such that $\xi\neq 0$ at some $x\in\pi^{-1}(U)$. We define a variation ${\tilde{g}}_{t}$ on $\pi^{-1}(U)$ by Theorem 3.6, setting $V_{n+1}=\xi$ and $V_{n+2}=\ldots=V_{n+p}=0$, and then use a non-zero, compactly supported in $U$ function $\rho$, such that $0\leq\rho\leq 1$ and $\rho(\pi(x))>0$, to define a variation $g_{t}$ on $M$ by $\partial_{t}g_{t}=(\rho\circ\pi)\cdot\partial_{t}{\tilde{g}}_{t}$. ∎
@@
To define a variation that preserves a Riemannian submersion with totally geodesic fibers, it is sufficient to assume that there exists a bounded Killing vector field $\xi$ on a fiber $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ for one $x\in N$, because all fibers of a Riemannian submersion with totally geodesic fibers are isometric and we can extend such $\xi$ to a vertical field on $(M,g_{0})$, whose restriction to every fiber is a Killing vector field. The isometries between totally geodesic fibers of $\pi$ on $(M,g_{0})$ are defined by flows of $g_{0}$-horizontal lifts of vector fields from $N$ [10].
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with minimal fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with minimal fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is divergence-free for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is divergence-free, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with totally umbilical fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally umbilical fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a conformal Killing field for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is conformal Killing, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
$2P_{{\cal H}}^{t}\partial_{t}(\nabla^{t}_{X}Y-\frac{1}{n}g_{t}(X,Y)H_{t})$ $=$ $\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{X}V_{i},Y)+g_{0}(\nabla^{0}_{Y}V_{i},X)$ (29)
$-\frac{2}{n}g_{t}(X,Y)\operatorname{div}^{0}_{{\cal V}}V_{i}\big{)}P_{{\cal H}}^{t}\pi^{*}W_{i}.$
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with minimal fibers. Then for every open set $U\subset M$ there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ such that $g_{t}=g_{0}$ on $M\setminus U$ and $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with minimal fibers for all $t\in(-\epsilon,\epsilon)$.
@@
Let $x\in\pi(U)$ and let ${\cal P}_{x}$ be the differential operator $\operatorname{div}^{0}_{{\cal V}}$ acting on the space $\mathfrak{X}_{{\cal F}_{x}}$ of vector fields on the fiber ${\cal F}_{x}$. Then ${\cal P}_{x}$ is underdetermined elliptic, i.e., its symbol is surjective and non-injective [2]. Hence, ${\cal P}_{x}$ has infinite-dimensional kernel ${\cal K}_{x}$, which includes an infinite set of compactly supported vector fields on ${\cal F}_{x}$ [6]. Let $\{V_{n+1},\ldots,V_{n+p}\}$ be such vector fields. We can assume that $\pi(U)$ is a normal coordinate neighbourhood of $x$; for every $y\in\pi(U)$ let $\phi_{x,y}:{\cal F}_{x}\rightarrow{\cal F}_{y}$ be the map induced by the flow of the $g_{0}$-horizontal lift of the vector field tangent to the geodesic from $x$ to $y$. Since the fibers are minimal, the maps $\phi_{x,y}$ preserve the volume form induced on the fibers (see e.g., [12, Chapter XV, Theorem 6.6]). Hence, extending $\{V_{n+1},\ldots,V_{n+p}\}$ to $U$ by maps $\phi_{x,y*}$, we obtain a set of vertical vector fields, such that the restriction to every fiber in $U$ of each of them is divergence-free and compactly supported on the fiber. Let $\rho$ be a smooth compactly supported function on $\pi(U)$. Multiplying each $V_{i}$ by $\rho\circ\pi$ we obtain a set of compactly supported, vertical vector fields in kernel of $\operatorname{div}^{0}_{{\cal V}}$, that by Theorem 3.6 defines a family $g_{t}$ with the required properties. ∎
@@ 4 Variational problems
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\operatorname{vol}_{g_{t}}$ be the volume form on $(M,g_{t})$. For $(0,2)$-tensor fields $Q,S$ on $M$, let
@@
where $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{{\cal V}}$ is a set of local $g_{0}$-orthonormal vector fields.
Let $J_{h},J_{H}:\operatorname{Riem}(M,{\cal V},g_{0})\rightarrow\mathbb{R}$ be functionals defined by formulas
@@
where $\|h_{t}\|^{2}=\sum\nolimits_{a=1}^{n}g_{t}(h_{t}(E_{a},E_{b}),h_{t}(E_{a},E_{b}))$ for a local $g_{t}$-orthonormal frame $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{{\cal V}}$. We say that a metric $g_{0}$ is critical for (30) (resp. (31)) with respect to variations preserving the Riemannian submersion $\pi$, if for all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$, such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$, we have $\partial_{t}J_{h}(g_{t})|_{t=0}=0$ (resp. $\partial_{t}J_{H}(g_{t})|_{t=0}=0$).
@@
We note that all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserving the Riemannian submersion $\pi$ also preserve the volume form $\operatorname{vol}_{g_{0}}$, since $\partial_{t}\operatorname{vol}_{g_{t}}=\frac{1}{2}\operatorname{Tr\,}B_{t}^{\sharp}\operatorname{vol}_{g_{t}}$ [3]. Indeed, let $\{E_{a},{\cal E}_{i}(t)\}$ be an adapted frame as in Lemma 3.8, then by (2) and (3) we have
@@ 4.1 Critical metrics
We will use notation from Section 2, especially its last paragraph. In particular, we have $P_{{\cal H}}^{0}(\delta_{0}h_{0})(X)=P_{{\cal H}}^{0}\sum\nolimits_{a=1}^{n}(\nabla^{0}_{E_{a}}h_{0})(E_{a},X)$ for all $X\in\mathfrak{X}_{{\cal V}}$ and any $g_{0}$-orthonormal local frame $\{E_{1},\ldots,E_{n}\}\subset\mathfrak{X}_{{\cal V}}$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Then $g_{0}$ is critical for (30), with respect to variations preserving the Riemannian submersion $\pi$, if and only if $P_{{\cal H}}^{0}(\delta_{0}h_{0})=0$.
@@
Let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a one-parameter family of Riemannian metrics on $M$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion for all $t\in(-\epsilon,\epsilon)$.
Let $\{U_{\alpha},\alpha\in\{1,\ldots,\Lambda\}\}$ be a finite open cover of $N$, such that on every $U_{\alpha}$ there exists a frame of linearly independent vector fields $\{W_{n+1}^{\alpha},\ldots,W_{n+p}^{\alpha}\}$ and on $\pi^{-1}(U_{\alpha})$ we have $\operatorname{vol}_{g_{0}}=\pi^{*}(\operatorname{vol}_{g_{N}}(x))\otimes\operatorname{vol}_{{\cal F}_{x},g_{0}}$ for all $x\in U_{\alpha}$ [12], where $\operatorname{vol}_{{\cal F}_{x},g_{0}}$ is the volume form of ${\cal F}_{x}$ induced from $(M,g_{0})$. Then, by Remark 3.7, there exist vertical vector fields $\{V_{n+1}^{\alpha}(t),\ldots,V_{n+p}^{\alpha}(t)\}$ on $\pi^{-1}(U_{\alpha})$ such that
$B_{t}^{\sharp}X=\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i}^{\alpha}(t),X)P_{{\cal H}}^{t}\pi^{*}W_{i}^{\alpha}$
for all vertical vector fields $X$ on $\pi^{-1}(U_{\alpha})$. Let $z\in\pi^{-1}(U_{\alpha})$, then for every set of local $g_{0}$-orthonormal vector fields $\{E_{a},a=1,\ldots,n\}\subset\mathfrak{X}_{{\cal V}}$ in a neighbourhood of $z$ we have, by (3.2), at the point $z$:
@@
$=2g_{t}((\partial_{t}\nabla^{t})_{E_{a}}E_{b},h_{t}(E_{a},E_{b}))=2g_{t}(P_{{\cal H}}^{t}(\partial_{t}\nabla^{t})_{E_{a}}E_{b},h_{t}(E_{a},E_{b}))$
$=\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{E_{a}}V^{\alpha}_{i}(t),E_{b})+g_{0}(\nabla^{0}_{E_{b}}V^{\alpha}_{i}(t),E_{a})\big{)}g_{t}(P_{{\cal H}}^{t}\pi^{*}W_{i}^{\alpha},h_{t}(E_{a},E_{b}))$
$=\sum\nolimits_{i=n+1}^{n+p}\langle\mathcal{L}_{V^{\alpha}_{i}(t)}g_{0},h^{i,\alpha}_{t}\rangle_{{\cal V}},$ (32)
@@
$=\sum\nolimits_{\alpha=1}^{\Lambda}\sum\nolimits_{i=n+1}^{n+p}\int_{\pi^{-1}(U_{\alpha})}f_{\alpha}\langle\mathcal{L}_{V_{i}^{\alpha}(0)}g_{0},h^{i,\alpha}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{g_{0}}$
$=\sum\nolimits_{\alpha=1}^{\Lambda}\sum\nolimits_{i=n+1}^{n+p}\int_{U_{\alpha}}\big{(}\int_{{\cal F}_{x}}\langle\mathcal{L}_{V_{i}^{\alpha}(0)}g_{0},h^{i,\alpha}_{0}\rangle_{{\cal V}}\,\operatorname{vol}_{{\cal F}_{x},g_{0}}\big{)}f_{\alpha}\operatorname{vol}_{g_{N}}(x).$ (33)
We will prove that $\partial_{t}J_{h}(g_{t})|_{t=0}=0$ for all variations $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserving the Riemannian submersion $\pi$ if and only if for all $x\in N$, all $V\in\mathfrak{X}_{{\cal F}_{x}}$ and all $W\in T_{x}N$ we have
@@
We then set $V_{n+2}=\ldots=V_{n+p}=0$ and define a variation ${\tilde{g}}_{t}$ on $\pi^{-1}(U)$ as in Theorem 3.6. Using a non-zero, compactly supported in $U$ function $\rho$, such that $0\leq\rho\leq 1$ we can define a variation $g_{t}$ on $M$ by $\partial_{t}g_{t}=\rho\cdot\partial_{t}{\tilde{g}}_{t}$, for which $\partial_{t}J_{h}(g_{t})|_{t=0}\neq 0$.
@@
We note that for all $X\in\mathfrak{X}_{{\cal V}}$ and for a local $g_{0}$-orthonormal adapted frame $\{E_{a},{\cal E}_{i}\}$, we can write the Codazzi formula [17] in the following form:
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. If the fibers of $\pi$ are totally umbilical on $(M,g_{0})$, then $g_{0}$ is critical for (30), with respect to variations preserving the Riemannian submersion $\pi$, if and only if all fibers of $\pi$ are of parallel mean curvature on $(M,g_{0})$, i.e., $P_{{\cal H}}^{0}\nabla^{0}_{X}H_{0}=0$ for all $X\in\mathfrak{X}_{{\cal V}}$.
@@
For totally umbilical fibers $P_{{\cal H}}^{0}(\delta_{0}h_{0})(X)=\frac{1}{n}P_{{\cal H}}^{0}\nabla^{0}_{X}H_{0}$ for all $X\in\mathfrak{X}_{{\cal V}}$. ∎
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion, where $M$ is compact. Then $g_{0}$ is critical for (31), with respect to variations preserving the Riemannian submersion $\pi$, if and only if $H_{0}$ is projectable.
@@
$\int_{{\cal F}_{x}}g_{0}(H_{0},\pi^{*}W)\cdot(\operatorname{div}^{0}_{{\cal V}}V)\;\operatorname{vol}_{{\cal F}_{x},g_{0}}=0.$ (35)
@@
It follows that $g_{0}$ is critical for (31) with respect to variations preserving the Riemannian submersion $\pi$ if and only if $P_{{\cal V}}^{0}\nabla^{0}g_{0}(H_{0},\pi^{*}W)=0$ for all $W\in\mathfrak{X}_{N}$. Writing $H_{0}=\sum\nolimits_{i=n+1}^{n+p}g_{0}(H_{0},\pi^{*}W_{i})\pi^{*}W_{i}$ for a local $g_{N}$-orthonormal basis $\{W_{n+1},\ldots,W_{n+p}\}$, we find that this condition is equivalent to $H_{0}$ being projectable, as for all $V\in\mathfrak{X}_{{\cal V}}$
@@
We note that for a fixed $x\in N$, we can define functionals by (30) and (31), with $M$ in their definitions replaced by the fiber ${\cal F}_{x}$, i.e.,
$J_{h,x}(g_{t})$ $=$ $\int_{{\cal F}_{x}}\|h_{t}\|^{2}\,\operatorname{vol}_{{\cal F}_{x},g_{t}},$ (36)
$J_{H,x}(g_{t})$ $=$ $\int_{{\cal F}_{x}}g_{t}(H_{t},H_{t})\,\operatorname{vol}_{{\cal F}_{x},g_{t}}.$ (37)
The critical metrics for (36) and (37) satisfy the same conditions as those given in Theorems 4.2 and 4.5, only on the fiber ${\cal F}_{x}$ rather than everywhere on $M$.